# Temperature Dependent Mechanical and Structural Properties of Uniaxially Strained Planar Graphene

### v1

@@ 3.3.1 Analytical expressions of bond length and bond angle distributions
Figure 10: Bond length distributions for selected temperatures $T$ (as shown in the legend), for stress along the armchair direction (left column) and the zigzag direction (right column), for increasing loads from top to bottom. In particular, the stresses $\sigma$ in the armchair direction are (a) 0.0813 eV/Å2, (b) 0.569 eV/Å2, (c) 1.06 eV/Å2, and (d) 1.55 eV/Å2. The stresses along the zigzag direction are (e) 0.0939 eV/Å2, (f) 0.751 eV/Å2, (g) 1.41 eV/Å2, and (h) 2.16 eV/Å2. The analytical expressions of Equation (22) are indicated by the solid curves.
@@
Figure 11: Bond angle distributions for different temperatures $T$ (as shown in the legend), for stress along the armchair direction (left column) and the zigzag direction (right column), for increasing values of stress from top to bottom. The stress in the armchair direction is (a) 0.0813 eV/Å2, (b) 0.569 eV/Å2, (c) 1.06 eV/Å2, and (d) 1.55 eV/Å2, while along the zigzag direction is (e) 0.0939 eV/Å2, (f) 0.751 eV/Å2, (g) 1.41 eV/Å2, and (h) 2.16 eV/Å2. The analytical expressions given by Equation (23) are indicated by the solid curves.
@@ 4 Conclusions
We investigated the planar dynamics of a graphene sheet using Hamiltonian formalism and an efficient symplectic integration technique allowing the creation of accurate numerical data for very long simulation times. Our MD simulations examined the effects of finite temperatures in the mechanical response of graphene. In particular, we derived stress-strain responses for two different directions of applied stress, along either the armchair or the zigzag direction, at various temperatures.
We obtained a small, almost linear decrease of the Young’s modulus of graphene as the temperature of the sheet increases. Such a variation of Young’s modulus with temperature is in line with previous investigations. Furthermore, an intriguing temperature dependence has been obtained for the third-order elastic modulus, which is found to decrease (slightly increase) its absolute value with increasing temperature, for stresses along the armchair (zigzag) direction. Finally, we found that tensile strength and failure strain decrease approximately linearly with temperature and quantified the slope of this variation.
We also discussed the dependence of the distributions of bond lengths and bond angles within the graphene sheet, on both the applied stress and temperature. Approximate analytical expressions for these distributions are provided. In particular, we found that the distributions can be described by the sum of two Gaussian peaks, where the center of each peak is obtained from the values of bond lengths or bond angles, respectively, in the strained graphene subjected to the particular amount of stress at zero temperature. The variance of each peak as a function of temperature can be derived by the corresponding data at zero applied stress, while for the bond length distributions a scaling factor is additionally incorporated to account for the highly anharmonic nature of the Morse potential. Thus, we have provided a detailed description of the effects of both stress and temperature on the structural properties of graphene.

### v4

@@ 3.3.1 Analytical expressions for the bond length and bond angle distributions
Figure 10: Bond length distributions at various temperatures $T$ (as shown in the legend) and applied stress along the armchair direction (left column) or the zigzag direction (right column), for increasing loads from top to bottom. In particular, the stresses $\sigma$ in the armchair direction are (a) 0.0813 eV/Å2, (b) 0.569 eV/Å2, (c) 1.06 eV/Å2, and (d) 1.55 eV/Å2. The stresses along the zigzag direction are (e) 0.0939 eV/Å2, (f) 0.751 eV/Å2, (g) 1.41 eV/Å2, and (h) 2.16 eV/Å2. The analytical expressions of Equation (22) are shown by solid curves and the corresponding numerical data by circles.
@@
Figure 11: Bond angle distributions for different temperatures $T$ (as shown in the legend) and stress along the armchair direction (left column) or the zigzag direction (right column), for increasing values of stress from top to bottom. The stress in the armchair direction is (a) 0.0813 eV/Å2, (b) 0.569 eV/Å2, (c) 1.06 eV/Å2, and (d) 1.55 eV/Å2, while along the zigzag direction is (e) 0.0939 eV/Å2, (f) 0.751 eV/Å2, (g) 1.41 eV/Å2, and (h) 2.16 eV/Å2. The analytical expressions given by Equation (23) are indicated by the solid curves and the corresponding numerical data by circles.
@@ 4 Conclusions
We investigated the planar dynamics of a uniaxially loaded graphene sheet using Hamiltonian formalism and an efficient symplectic integration technique allowing the creation of accurate numerical data for very long simulation times. Our MD simulations examined the effects of thermal fluctuations in the mechanical response of graphene. In particular, we derived stress-strain responses for two different directions of applied stress, along either the armchair or the zigzag direction, at various temperatures.
A small, almost linear decrease of the Young modulus of graphene is obtained as the temperature of the sheet increases. Such a variation of Young modulus with temperature is in line with previous investigations. Furthermore, an intriguing temperature dependence of the third-order elastic modulus has been presented for the first time, which is found to decrease (slightly increase) its absolute value with increasing temperature, for stresses along the armchair (zigzag) direction. Finally, we found that the tensile strength and failure strain decrease approximately linearly with temperature and computed the slope of this variation. A quantitative comparison with existing results regarding these variations is presented.
It is worth mentioning that even though our model is restricted to planar deformations, results obtained for the intrinsic strength and fracture strain at room temperature are in agreement with MD simulations ref-Zhao-NanoLet-2009 ; ref-Varillas-MechSci-2024 ; ref-Lee-NanoResLets-2015 allowing out-of-plane displacements of carbon atoms. Moreover, the values of Young modulus and fracture strength at 300 K are in accordance with experimental estimates presented in References ref-Lee-Science-2008 ; ref-Cao-NatComm-2020 ; ref-Varillas-MechSci-2024 and Reference ref-Lee-Science-2008 , respectively.
The dependence of the distributions of bond lengths and bond angles within the graphene sheet, on both the applied stress and temperature is also discussed. Approximate analytical expressions for these distributions are provided. In particular, we found that the distributions can be described by the sum of two Gaussian peaks, where the center of each peak is obtained from the values of bond lengths or bond angles, respectively, in the strained graphene subjected to the particular amount of stress at zero temperature. The variance of each peak as a function of temperature can be derived by the corresponding data at zero applied stress, while for the bond length distributions a scaling factor is additionally incorporated to account for the anharmonicity of the Morse potential. Thus, for the first time, a detailed description of the effects of both stress and temperature on the structural properties of graphene is reported.
@@
Author contributions: Conceptualization, Ch. S. and G. K.; methodology, Ch. S. and G. K.; software, S. E.; validation, S. E., Ch. S. and G. K.; formal analysis, S. E., Ch. S. and G. K.; investigation, S. E., Ch. S. and G. K.; writing—review and editing, S. E., Ch. S. and G. K.; visualization, S. E.; supervision, Ch. S. and G. K. All authors have read and agreed to the published version of the manuscript.